Solve each system by substitution method. Indicate whether each system is independent, inconsistent or dependent.
3y=y+5
3x-9y= -10

and
x=1/8y-1
y=1/4x +39

Answer by venugopalramana(3286)

Solve each system by substitution method. Indicate whether each system is independent, inconsistent or dependent.
3y=y+5...............IS THIS 3X=Y+5?CHECKUP..ASSUMING SO.PUTTING THIS IN THE EQN.
3x-9y= -10
Y+5-9Y=-10
-8Y=-10-5=-15
Y=15/8
3X=15/8 + 5 = (15+40)/8 =55/8
X=55/24
SYSTEM IS INDEPENDENT